# Two-stage Generative Question Answering on Temporal Knowledge Graph Using Large Language Models

### v1

@@ Abstract
Temporal knowledge graph question answering (TKGQA) poses a significant challenge task, due to the temporal constraints hidden in questions and the answers sought from dynamic structured knowledge. Although large language models (LLMs) have made considerable progress in their reasoning ability over structured data, their application to the TKGQA task is a relatively unexplored area. This paper first proposes a novel generative temporal knowledge graph question answering framework, GenTKGQA, which guides LLMs to answer temporal questions through two phases: Subgraph Retrieval and Answer Generation. First, we exploit LLM’s intrinsic knowledge to mine temporal constraints and structural links in the questions without extra training, thus narrowing down the subgraph search space in both temporal and structural dimensions. Next, we design virtual knowledge indicators to fuse the graph neural network signals of the subgraph and the text representations of the LLM in a non-shallow way, which helps the open-source LLM deeply understand the temporal order and structural dependencies among the retrieved facts through instruction tuning. Experimental results demonstrate that our model outperforms state-of-the-art baselines, even achieving 100% on the metrics for the simple question type.
Two-stage Generative Question Answering on Temporal Knowledge
Graph Using Large Language Models
[…]
@@ 1 Introduction
Real-world knowledge is frequently updated rather than static Erxleben et al. (2014); Boschee et al. (2015), e.g., (Obama, hold_position, President) is merely valid only for a certain period [2009, 2016]. Hence, the temporal knowledge graph (TKG) is proposed as a database for storing dynamic structured facts associated with timestamps, denoted as (subject, relation, object, timestamp). Temporal knowledge graph question answering (TKGQA) aims to answer a natural question with explicit or implicit temporal constraints based on the TKG, e.g., "Who held the position of president (in 2017) or (after Obama)?". Due to the temporal constraints hidden in questions and the answers being sought from dynamic structured knowledge, TKGQA is one of the most challenging QA tasks.
@@
Recently, large language models (LLMs) have shown strong competitiveness in various fields Fei et al. (2023); Ye et al. (2023). Some researchers explore the reasoning capability of LLMs for structured knowledge based on KGQA tasks Baek et al. (2023); Kim et al. (2023), and some works examine the temporal reasoning capabilities of LLMs through time-sensitive QA tasks Chen et al. (2021a); Tan et al. (2023). Intuitively, LLMs have the ability to deal with temporal structured knowledge. Based on the above findings, we attempt to utilize LLMs for the TKGQA task and summarize the following two challenges: 1) Question-relevant Subgraph Retrieval. A common practice to enhance the LLM’s domain-specific reasoning capability is to input query-relevant information as additional knowledge into the LLMs Hu et al. (2023). As shown in Figure 1, ChatGPT OpenAI (2023) cannot answer temporal questions directly, but it can answer simple questions when the relevant context is provided. However, finding facts relevant to the problem is a struggle due to the large search space with both structural and temporal dimensions. For example, for the complex temporal question "Who held the position of president after Obama?", the structural link hold position between the entities and the temporal constraints after 2016 in the problem are unknown, so directly finding all relevant facts about the given entities Obama and President is bound to introduce too much noisy information. How to accurately retrieve relevant facts from a two-dimensional space is the first challenge. 2) Complex-type Question Reasoning. Recent research about LLM-based KG reasoning mostly inputs structured knowledge in the natural text form into the task prompt and reasons about the answers in a training-free method Yang et al. (2024). However, this approach fuses subgraph information with the LLM in a shallow manner, which limits the inference performance on the complex question type. As illustrated in Figure 1, ChatGPT is unable to understand the chronological order of the relevant facts and answers incorrectly on the complex question type "after". How subgraph information can be integrated into LLM representations in a non-superficial way to simulate structured reasoning remains an open question.
@@
Hence, we propose GenTKGQA, a novel generative temporal knowledge graph question answering framework consisting of two phases, subgraph retrieval and answer generation, which is used to address the above two challenges, respectively. At the first phase, we find that the structural and temporal scope of the subgraph is determined by the relation links and the temporal constraints in the question, respectively. Therefore, we use a divide-and-conquer strategy to reduce the subgraph search space by decomposing the complex subgraph retrieval problem into two subtasks, namely, relation ranking and time mining. Then, we utilize the LLM’s internal knowledge to mine structural connections between entities and time constraints in the problem without extra training. We only need to input few-shot examples into the prompt to accomplish subgraph retrieval of the entire data. At the second phase, we fine-tune the open-source LLM with instruction tuning to incorporate structural and temporal information of the subgraph in a non-shallow way. Recent works illustrate that fusing graph neural network (GNN) representations and language text representations can enhance the ability of LMs to perceive graph structure Zhang et al. (2022b). Thus, we design three novel virtual knowledge indicators to bridge the links between pre-trained GNN signals of the temporal subgraph and text representations of the LLMs, which guides the LLMs in deeply understanding the graph structure and improves their reasoning ability for complex temporal questions. Overall, our contribution can be summarized in the following four points:
@@
2) We motivate the LLM’s intrinsic knowledge to mine the temporal constraints and structural connections in the questions without extra training, which reduces the subgraph search space from both structure and time dimensions.
@@
4) Experiment results show that GenTKGQA as a generative QA model performs consistently better than embedding-based QA methods, even reaching 100% on all metrics for the simple question type.
@@ 2.1 TKGQA Methods
Temporal knowledge graph question answering (TKGQA) task aims to answer complex questions in the natural language format using entities and timestamps from the given TKG Jia et al. (2018, 2021); Saxena et al. (2021); Chen et al. (2023). Existing mainstream methods employ TKG embeddings to represent the entities, relations and timestamps, and use the TKGE scoring function to select the entity or time with the highest relevance as the answer Saxena et al. (2021). However, single embedding methods have difficulty handling complex reasoning problems with implicit time constraints. Therefore, recent methods try to incorporate other modules to improve the model performance on complex problems. Specifically, TSQA Chen et al. (2021b) proposes a contrastive approach to enhance the time sensitivity of the model. TempoQR Mavromatis et al. (2022) designs three modules, namely context, entity, and time-aware information, to enhance the incorporation of the TKG into questions. LGQA Liu et al. (2023b) applies a multi-hop message passing GNN layer to combine the global and local information of the given problem. Despite the effectiveness of these approaches, few works explore how approaches other than the embedded-based method can address the TKGQA task.
@@ 2.2 LMs for Temporal Question Answering
Language models (LMs) have exhibited strong performance on the question answering task Raffel et al. (2020). In recent years, some researchers have explored the temporal reasoning capabilities of LMs and propose several typical time-sensitive QA datasets. They focus on temporal question answering either within a closed-book setting to assess models’ internal memorization of temporal facts Liska et al. (2022); Dhingra et al. (2022), or within an open-book setting to evaluate models’ temporal understanding and reasoning capability over unstructured texts Zhang and Choi (2021); Chen et al. (2021a); Tan et al. (2023). In the context of the latter setting, some works propose to use the graph structure extracted from text to assist the model in determining the temporal order between events Mathur et al. (2022); Su et al. (2023); Yang et al. (2023), which is similar but fundamentally different from our work. These approaches aim to answer temporal questions based on the known natural text context. In contrast, our model focuses on structured temporal knowledge as auxiliary information that needs to be retrieved by the model.
@@ 2.3 LMs for KG Question Answering
How to combine LMs and KG for question answering has become a hot issue. Some works attempt to enhance question representation and relation matching with PLMs in the multi-hop KGQA task Saxena et al. (2020); Zhang et al. (2022a); Jiang et al. (2023b), but there is no interaction between the LM and KG representations. Other works try to use one modality to ground the other, i.e., using the encoded representation of a linked KG to augment the text representation Lin et al. (2019); Yang et al. (2019), or using the text representation of the PLM to enhance the graph reasoning model Feng et al. (2020a). The most recent approaches enable deep integration of the two modalities by jointly updating the GNN and LM representation Yasunaga et al. (2021); Zhang et al. (2022b).
However, the emergence of large language models (LLMs) has changed how LMs handle the KGQA task, which is divided into two main approaches: training-free and fine-tuning Yang et al. (2024). Recent works attempt to append query-relevant facts as the input prompt for LLMs and make inferences without extra training Baek et al. (2023); Jiang et al. (2023a); Kim et al. (2023); Li et al. (2024). Fine-tuning the full parameters of the LLM can be cost-prohibitive. Hence, KPE Zhao et al. (2023) enables knowledge integration by freezing PLM parameters and introducing trainable parameter adapters. ChatKBQA Luo et al. (2023) employs the LoRA Hu et al. (2022) technique to fine-tune open-source LLMs, achieving the logical query form generation. Applying LLMs to the temporal KGQA task remains an unexplored area.
@@ 3 Preliminaries
A temporal knowledge graph (TKG) $\mathcal{G}:=(\mathcal{E},\mathcal{R},\mathcal{T},\mathcal{F})$ is a multi-relational, directed graph with timestamped edges between entities, where $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ represent the sets of entities, relationships and timestamps, respectively. Each fact in the $\mathcal{G}$ can be represented as a quadruple $(s,r,o,t)\in\mathcal{F}$, corresponding to entity $s/o\in\mathcal{E}$, relation type $r\in\mathcal{R}$ and timestamp $t\in\mathcal{T}$. Given a natural language question $q$, TKGQA aims to extract an entity $s/o$ or a timestamp $t$ that correctly answers the question.
@@ 4 Method
Figure 2: The overall architecture of our proposed GenTKGQA can be divided into two stages, subgraph retrieval and answer generation. Given a temporal problem, we mark the entities provided, the implied time constraints and the links between the entities with underline, blue font and orange font, respectively.
@@ 4.1 Subgraph Retrieval
We split the complex subgraph retrieval problem into a relation ranking sub-task in the structural dimension and a time mining sub-task in the temporal dimension. With such a strategy, we only need to provide a small number of samples to complete the subgraph retrieval.
@@ 4.1.2 Time Mining
We find that natural questions contain temporal constraints, either explicit or implicit, and we can easily determine the range of relevant facts in the temporal dimension by using explicit temporal constraints such as "in 2008", "at the year of 2012", etc. How to capture implicit temporal constraints is the key to improving the efficiency of searching for relevant facts. For example, for the complex temporal question "Who held the position of president after Obama?", the implicit temporal restriction is known to be (after 2016) based on the temporal validity (2009, 2016) of the fact (Obama, hold_position, President). We design specific prompt templates based on different answer types as well as consider the temporal validity of the question-containing facts (given entities and relations matched in the previous section) to get the temporal constraints. The details of the templates are shown in Appendix C.
@@ 4.2 Answer Generation
In this section, we will discuss how to incorporate the knowledge retrieved in the previous section 4.1 into the LLM. Previous fundamental approaches to incorporate KG structural information focus on adding the knowledge to the input prompt in the text form, e.g., (subject, relation, object). However, incorporating query-relevant facts into LLMs in text form is not a good choice. The reason is that this shallow interaction does not enable the model to understand the structural dependencies and temporal order between facts, leading to weak temporal reasoning in the complex problem type.
@@ 4.2.1 Temporal GNN
where $\mathbf{e}_{i}^{(l-1)}$ is the entity representation of $e_{i}$ at the $l$-1 layer, $\mathbf{r}_{ij}$ and $\mathbf{t}_{ij}$ are the embeddings of the relation and timestamp connecting $e_{i}$ and $e_{j}$. $\mathbf{W}_{m}$ is a linear transformation. Next, the node representation $\mathbf{e}_{j}^{(l)}$ is calculated via message passing between neighbors on the $\mathcal{G}_{sub}$:
@@
$\mathbf{W}_{q},\mathbf{W}_{k}$ are linear transformations, $f_{n}$ is the RELU activation function. Through the above process, we obtain entity representations $\mathbf{e}_{j}$ with subgraph structural and temporal information. Following embedding-based TKGQA methods, we use the link prediction task to pre-train the graph neural network representations. Specifically, for each fact $(s,r,o,t)$ in the TKG, we generate a query $(s,r,[mask],t)$ or $([mask],r,o,t)$ by masking the object or subject entity. Then, we obtain the mask embedding $\mathbf{e}_{[mask]}^{(l)}$ through Eq. (2), and feed it into the multi-layer perceptron (MLP) decoder to maximize the probability of the missing entity $o$ and $s$ through the cross-entropy loss function:
@@ 4.2.2 Virtual Knowledge Indicators
where $\mathbf{t}_{min}$ and $\mathbf{t}_{max}$ denote the embeddings for the minimu and maximum values of time in the subgraph, respectively. The intuition follows BERT that use position embeddings for tokens Devlin et al. (2019). Here, time embeddings can be seen as entity positions in the time dimension. The relation and object indicators are the same as subject. At last, we employ a simple linear layer $\mathbf{W}_{p}$ to project them into the textual representation space of the LLM. The final input prompt sequence $\mathcal{S}=\mathcal{V}:\mathcal{I}:\mathcal{Q}:\mathcal{A}$, $\mathcal{V}$ represent virtual indicator tokens. Details of the instruction template can be found in the Appendix C. The optimization objective of the LLM $\mathcal{M}$ can be formulated as:

### v2

@@ Abstract
Temporal knowledge graph question answering (TKGQA) poses a significant challenge task, due to the temporal constraints hidden in questions and the answers sought from dynamic structured knowledge. Although large language models (LLMs) have made considerable progress in their reasoning ability over structured data, their application to the TKGQA task is a relatively unexplored area. This paper first proposes a novel generative temporal knowledge graph question answering framework, GenTKGQA, which guides LLMs to answer temporal questions through two phases: Subgraph Retrieval and Answer Generation. First, we exploit LLM’s intrinsic knowledge to mine temporal constraints and structural links in the questions without extra training, thus narrowing down the subgraph search space in both temporal and structural dimensions. Next, we design virtual knowledge indicators to fuse the graph neural network signals of the subgraph and the text representations of the LLM in a non-shallow way, which helps the open-source LLM deeply understand the temporal order and structural dependencies among the retrieved facts through instruction tuning. Experimental results on two widely used datasets demonstrate the superiority of our model.
Two-stage Generative Question Answering on Temporal Knowledge Graph Using Large Language Models
Yifu Gao1, Linbo Qiao1*, Zhigang Kan1, Zhihua Wen1, Yongquan He3, Dongsheng Li1,2* 1 National Key Laboratory of Parallel and Distributed Computing, National University of Defense Technology, Changsha, China 2 Xiangjiang Laboratory, Changsha, China 3 Meituan, Beijing, China {gaoyifu, qiao.linbo, kanzhigang13, zhwen, dsli}@nudt.edu.cn heyongquan@meituan.com
*Corresponding Author
@@ 1 Introduction
Real-world knowledge is frequently updated rather than static Erxleben et al. (2014); Boschee et al. (2015), e.g., (Obama, hold_position, President) is merely valid only for a certain period [2009, 2016]. Hence, the temporal knowledge graph (TKG) is proposed as a database for storing dynamic structured facts associated with timestamps, denoted as (subject, relation, object, timestamp). Temporal knowledge graph question answering (TKGQA) aims to answer a natural question with explicit or implicit temporal constraints based on the TKG, e.g., "Who held the position of president (in 2017) or (after Obama)?". Due to the temporal constraints hidden in questions and the answers sought from dynamic structured knowledge, TKGQA is one of the most challenging QA tasks.
@@
Recently, large language models (LLMs) have shown strong competitiveness in various fields Fei et al. (2023); Ye et al. (2023). Some researchers explore the reasoning capability of LLMs for structured knowledge based on KGQA tasks Baek et al. (2023); Kim et al. (2023), and some works examine the temporal reasoning capabilities of LLMs through time-sensitive QA tasks Chen et al. (2021); Tan et al. (2023). Intuitively, LLMs have the ability to deal with temporal structured knowledge. Based on the above findings, we attempt to utilize LLMs for the TKGQA task and summarize the following two challenges: 1) Question-relevant Subgraph Retrieval. A common practice to enhance the LLM’s domain-specific reasoning capability is to input query-relevant information as additional knowledge into the LLMs Hu et al. (2023). As shown in Figure 1, ChatGPT  [^1] https://openai.com/blog/chatgpt cannot answer temporal questions directly, but it can answer simple questions when the relevant facts are provided. However, finding facts relevant to the problem is a struggle due to the large search space with both structural and temporal dimensions. For example, for the complex temporal question "Who held the position of president after Obama?", the structural link hold position between the entities and the temporal constraints after 2016 in the problem are unknown, so directly finding all relevant facts about the given entities Obama and President is bound to introduce too much noise information. How to accurately retrieve relevant facts from a two-dimensional space is the first challenge. 2) Complex-type Question Reasoning. Recent works about LLM-based KG reasoning mostly input structured knowledge in the natural text form into the task prompt and reason about the answers in a training-free manner Yang et al. (2024). However, these approaches fuse subgraph information with the LLM in a shallow way, which limits the inference performance on the complex question type. As illustrated in Figure 1, ChatGPT cannot understand the chronological order of the relevant facts and answers incorrectly on the complex question type "final". How subgraph information can be integrated into LLM representations in a non-superficial way to simulate structured reasoning remains an open question.
@@
Hence, we propose GenTKGQA, a novel generative temporal knowledge graph question answering framework consisting of two phases, subgraph retrieval and answer generation, which is used to address the above two challenges, respectively. At the first phase, we find that the structural and temporal scope of the subgraph is determined by the relation links and the temporal constraints in the question, respectively. Therefore, we use a divide-and-conquer strategy to reduce the subgraph search space by decomposing the complex subgraph retrieval problem into two subtasks, namely, relation ranking and time mining. Then, we utilize the LLM’s internal knowledge to mine structural connections between entities and time constraints in the problem without extra training. We only need to input few-shot examples into the prompt to accomplish subgraph retrieval of the entire data. At the second phase, we fine-tune the open-source LLM with instruction tuning to incorporate structural and temporal information of the subgraph in a non-shallow way. Recent works illustrate that fusing graph neural network (GNN) representations and language text representations can enhance the ability of LMs to perceive graph structure Zhang et al. (2022b). Thus, we design three novel virtual knowledge indicators to bridge the links between pre-trained GNN signals of the temporal subgraph and text representations of the LLM, which guides the LLMs in deeply understanding the graph structure and improves their reasoning ability for complex temporal questions. Overall, our contribution can be summarized in the following four points:
@@
2) We motivate the LLM’s intrinsic knowledge to mine the temporal constraints and structural connections in the questions without extra training, which reduces the subgraph search space from both structural and temporal dimensions.
@@
4) Experiment results on two widely used datasets show that GenTKGQA as a generative model performs consistently better than embedding-based methods on the Hits@1 metric.
@@ 2.1 TKGQA Methods
Temporal knowledge graph question answering (TKGQA) task aims to answer complex questions in the natural language format using entities and timestamps from the given TKG Jia et al. (2018, 2021); Chen et al. (2023b). Existing mainstream methods employ TKG embeddings to represent the entities, relations and timestamps, and use the scoring function to select the entity or time with the highest relevance as the answer Saxena et al. (2021). However, single embedding methods have difficulty handling complex reasoning problems with implicit time constraints. Therefore, recent methods try to incorporate other modules to improve the model performance on complex problems. Specifically, TSQA Shang et al. (2022) proposes a contrastive approach to enhance the model’s time sensitivity. TempoQR Mavromatis et al. (2022) designs three modules, namely context, entity and time-aware information, to enhance the incorporation of the TKG into questions. Besides, some approaches propose to solve the TKGQA task with problem-relevant subgraph reasoning Chen et al. (2022); Liu et al. (2023b); Sharma et al. (2023). Despite the effectiveness of these approaches, few studies have explored how LLMs can address the TKGQA task.
@@ 2.2 LMs for Temporal Question Answering
Language models (LMs) have exhibited strong performance on the question answering task Raffel et al. (2020). In recent years, some researchers have explored the temporal reasoning capabilities of LMs and propose several typical time-sensitive QA datasets. They focus on temporal question answering either within a closed-book setting to assess models’ internal memorization of temporal facts Liska et al. (2022); Dhingra et al. (2022), or within an open-book setting to evaluate models’ temporal understanding and reasoning capability over unstructured texts Zhang and Choi (2021); Chen et al. (2021); Tan et al. (2023). In the context of the latter setting, some works propose to use the graph structure extracted from text to assist the model in determining the temporal order between events Mathur et al. (2022); Su et al. (2023); Yang et al. (2023); Xiong et al. (2024), which is similar but fundamentally different from our work. These approaches aim to answer temporal questions based on the known natural text context. In contrast, our model focuses on structured temporal knowledge as auxiliary information that needs to be retrieved by the model.
@@ 2.3 LMs for KG Question Answering
How to combine LMs and KG for question answering has become a hot issue. Some works attempt to enhance question representation and relation matching with PLMs in the multi-hop KGQA task Saxena et al. (2020); Zhang et al. (2022a); Jiang et al. (2023b), but there is no interaction between the LM and KG representations. Other works try to use one modality to ground the other, i.e., using the encoded representation of a linked KG to augment the text representation Lin et al. (2019); Yang et al. (2019), or using the text representation of the PLM to enhance the graph reasoning model Feng et al. (2020a). The most recent approaches enable deep integration of the two modalities by jointly updating the GNN and LM representations Yasunaga et al. (2021); Zhang et al. (2022b).
However, the emergence of large language models (LLMs) has changed how LMs handle the KGQA task, which is divided into two main approaches: training-free and fine-tuning Yang et al. (2024). Recent works attempt to append query-relevant facts as the input prompt for LLMs and make inferences without extra training Baek et al. (2023); Wu et al. (2023); Jiang et al. (2023a); Kim et al. (2023); Li et al. (2024). Fine-tuning the full parameters of the LLM can be cost-prohibitive. Hence, KPE Zhao et al. (2023) enables knowledge integration by freezing PLM parameters and introducing trainable parameter adapters. ChatKBQA Luo et al. (2023) employs the LoRA Hu et al. (2022) technique to fine-tune open-source LLMs, achieving the logical query form generation. Besides, KoPA Zhang et al. (2023) incorporates the KG embeddings into the LLMs with a prefix adapter, aiming to achieve structrual-aware reasoning in the LLMs. Applying LLMs to the temporal KGQA task remains an unexplored area.
@@ 3 Preliminaries
A temporal knowledge graph (TKG) $\mathcal{G}:=(\mathcal{E},\mathcal{R},\mathcal{T},\mathcal{F})$ is a multi-relational, directed graph with timestamped edges between entities, where $\mathcal{E}$, $\mathcal{R}$ and $\mathcal{T}$ represent the sets of entities, relations and timestamps, respectively. Each fact in the $\mathcal{G}$ can be represented as a quadruple $(s,r,o,t)\in\mathcal{F}$, corresponding to entity $s/o\in\mathcal{E}$, relation type $r\in\mathcal{R}$ and timestamp $t\in\mathcal{T}$. Given a natural language question $q$, TKGQA aims to extract entities $s/o$ or timestamps $t$ that correctly answer the question $q$.
@@ 4 Method
Figure 2: The overall architecture of our proposed GenTKGQA can be divided into two stages, subgraph retrieval and answer generation. Given a temporal question, we mark the entities provided, the implied time constraints and the links between the entities with underline, blue font and orange font, respectively.
@@ 4.1 Subgraph Retrieval
We split the complex subgraph retrieval problem into a relation ranking subtask in the structural dimension and a time mining subtask in the temporal dimension. With such a strategy, we only need to provide a small number of samples to complete the subgraph retrieval.
@@ 4.1.2 Time Mining
We find that natural questions contain temporal constraints, either explicit or implicit, and we can easily determine the range of relevant facts in the temporal dimension by using explicit temporal constraints such as "in 2008", "at the year of 2012", etc. How to capture implicit temporal constraints is the key to improving the efficiency of searching for relevant facts. For example, for the complex temporal question "Who held the position of president after Obama?", the implicit temporal restriction is known to be (after 2016) based on the temporal validity (2009, 2016) of the fact (Obama, hold_position, President). We design specific prompt templates based on different answer types as well as consider the temporal validity of the facts in the questions (given entities and relations matched in the Section 4.1.1) to get the temporal constraints. The details of the templates are shown in Appendix C.
@@ 4.2 Answer Generation
In this section, we will discuss how to incorporate the knowledge retrieved in the previous section 4.1 into the LLM. Previous fundamental approaches to incorporate KG structural information focus on adding the knowledge to the input prompt in the text form, i.e., (subject, relation, object). However, incorporating query-relevant facts into LLMs in text form is not a good choice. Because such shallow interactions do not enable the model to understand the structural dependencies and temporal order between facts, leading to weak temporal reasoning in the complex problem type.
@@ 4.2.1 Temporal GNN
where $\mathbf{e}_{i}^{(l-1)}$ is the entity representation of $e_{i}$ at the $l$-1 layer, $\mathbf{r}_{ij}$ and $\mathbf{t}_{ij}$ are the embeddings of the relation and timestamp connecting $e_{i}$ and $e_{j}$. $\mathbf{W}_{m}$ is a linear transformation. Next, the node representation $\mathbf{e}_{j}^{(l)}$ is calculated via message passing between neighbors on the $\mathcal{G}_{sub,z}$:
@@
$\mathbf{W}_{q},\mathbf{W}_{k}$ are linear transformations, $f_{n}$ is the RELU activation function. Through the above process, we obtain the entity representation $\mathbf{e}_{j}$ with subgraph structural and temporal information. Following embedding-based TKGQA methods, we use the link prediction task to pre-train the graph neural network representations. Specifically, for each fact $(s,r,o,t)$ in the TKG, we generate a query $(s,r,[mask],t)$ or $([mask],r,o,t)$ by masking the object or subject entity. Then, we obtain the mask embedding $\mathbf{e}_{[mask]}^{(l)}$ through Eq.(2), and feed it into the multi-layer perceptron (MLP) decoder to maximize the probability of the missing entity $o$ and $s$ through the cross-entropy loss function:
@@ 4.2.2 Virtual Knowledge Indicators
where $\mathbf{t}_{min}$ and $\mathbf{t}_{max}$ denote the embeddings for the minimum and maximum values of time in the subgraph, respectively. The intuition follows BERT that use position embeddings for tokens Devlin et al. (2019). Here, time embeddings can be seen as entity positions in the time dimension. The relation and object indicators are the same as subject. At last, we employ a simple linear layer $\mathbf{W}_{p}$ to project them into the textual representation space of the LLM. The final input prompt sequence $\mathcal{S}=\mathcal{V}:\mathcal{I}:\mathcal{Q}:\mathcal{A}$, $\mathcal{V}$ represent virtual indicator tokens. Details of the instruction template can be found in the Appendix C. The optimization objective of the LLM $\mathcal{M}$ can be formulated as: